OpenStudy (anonymous):

Find g(f(x)) when f(x)= x^2+1 and g(x)= sqrt(x+3)

OpenStudy (anonymous):

\[g\circ f(x)=g(f(x))=g(x^2+1)=\sqrt{(x^2+1)+3}=\sqrt{x^2+4}\]

OpenStudy (anonymous):

\[\sqrt{x^2+4}\] would be my final answer, correct? =) @satellite73